A pack contains n cards numbered 1 to n.Two consecutive numbered cards are removed from the pack and the sum of the remaining cards is 1224.If the smaller of the numbers on the removed cards is k,  then k-20=

Grade:11

1 Answers

Arun
5 years ago
Acc to given condition n
n(n+1)/2 -(k)-(k+1)=1224
get k= n(n+1)-2450 / 4
since n should be such that n(n+1)>2450 and k is integer only n= 50 satisfies , 
then k= 25 and k-20 = 5